# CuLEmo: Cultural Lenses on Emotion - Benchmarking LLMs for Cross-Cultural Emotion Understanding

### v2

@@ Abstract
Tadesse Destaw Belay1,3, Ahmed Haj Ahmed2, Alvin Grissom II2, Iqra Ameer4, Grigori Sidorov1, Olga Kolesnikova1, Seid Muhie Yimam5 1Instituto Politécnico Nacional, 2Haverford College, 3Wollo University, 4Pennsylvania State University, 5University of Hamburg, Contact: tadesseit@gmail.com
@@ 1 Introduction
LLMs and agent systems are employed to interact extensively with humans across applications such as customer service, healthcare, and education Wang et al. (2024). To facilitate effective interaction, incorporating aspects of cognitive and emotional-social intelligence—the ability to recognize and interpret human emotions Mathur et al. (2024)—can facilitate better interactions with more people. Factors such as age, cultural background, and personal experiences influence how individuals perceive and process information, particularly within subjective NLP tasks. Among these, emotion recognition (ER) and sentiment analysis (SA) are particularly sensitive to language- and culture-specific nuances Plaza-del-Arco et al. (2024).
@@ 3.1 Collecting Cultural Events
We manually craft scenarios, search on the web, and prompt LLMs to gather traditions, events, norms, and actions that elicit culturally different emotions across six target countries (UAE, USA, Germany, Ethiopia, India, and Mexico). We draw inspiration from the work of Havaldar et al. (2023b) to enhance topic diversity. Language representatives are asked to propose events distinct from those of other countries in the form of emotion-oriented questions. Importantly, these events do not contain explicit emotional keywords (as typically seen in traditional emotion datasets). We also refer to the International Survey on Emotion Antecedents and Reactions (ISEAR) (Scherer and Wallbott, 1994) data format, "When I … situations that cause a specific emotion"—a well-known English dataset for emotion analysis consisting of self-reported events from around 3,000 respondents across 37 countries and five continents. Table 1 displays the ten broad categories of the CuLEmo dataset and the number of questions in each category.
@@
Table 1: Ten broad categories and the number of events (# of Qn - number of questions) in the CuLEmo dataset.
@@ 3.2 Human-Adapted Translation
After collecting the events in English, we translate them into five target languages—Arabic, Amharic, German, Hindi, and Spanish—using Google Translate, followed by native-speaker approvals. Because the questions are simple “How do you feel when … ?” questions and lack explicit emotion keywords, translation quality did not affect their cultural content. While we acknowledge that existing works often depend on translating emotion-annotated data from English into other languages with their labels Kajava et al. (2020); Tahir et al. (2023); Bianchi et al. (2022), our translation process is done before any annotation. To ensure correctness, native speakers evaluate the translations. While most translations were acceptable, a few adjustments were made—e.g., to fix gender references and timing expressions for Amharic.
@@ 3.4 CuLEmo Annotation
We use Amazon Mechanical Turk (MTurk) for most of our annotations, ensuring at least five native-speaker annotators per instance from the respective targeted country. We use a customized POTATO Pei et al. (2022) annotation tool for languages lacking sufficient MTurk annotators (e.g., Amharic) and recruit local native speakers who met our criteria. Annotators are fairly compensated \$12/hr. (Prolific [^3] https://www.prolific.com/’s minimum annotation wage is \$9/hr). In cases where no majority vote emerges among five annotators, we assign two additional annotations and use that majority vote.
@@ 5.1 Culture-Aware Emotion Prediction
Results: Table 2 presents the accuracy of each LLM for culture-aware emotion prediction. The choice of prompt language significantly influences performance. Generally, proprietary models are less affected by in-language prompts compared to open-source models, especially for Spanish and German. Certain cultures appear better represented in the models—Ministral-8B scores highly on German (72%), and GPT-4 performs best on Mexican (65%). In contrast, performance in Indian culture (Hindi) lags, particularly those using Hindi, Amharic, or Arabic scripts. Larger models do not always outperform smaller ones; Gemma-2-2B and Ministral-8B show competitive or superior accuracy relative to some proprietary models. When prompted in English, all models achieve a reasonable accuracy. Ministral-8B can exceed proprietary performance in English and German. GPT-4 answers the same emotion while using the corresponding country name in the prompting; see predicted examples in Table 3. Overall, results suggest that culture-aware emotion understanding remains challenging for the tested LLMs, especially for lesser-resourced languages and cultures.

### v3

@@ Abstract
Tadesse Destaw Belay1,3, Ahmed Haj Ahmed2, Alvin Grissom II2, Iqra Ameer4, Grigori Sidorov1, Olga Kolesnikova1, Seid Muhie Yimam5 1Instituto Politécnico Nacional, 2Haverford College, 3Wollo University, 4Pennsylvania State University, 5University of Hamburg Correspondence: tadesseit@gmail.com
@@ 1 Introduction
LLMs and agent systems are employed to interact extensively with humans across applications such as customer service, healthcare, and education Wang et al. (2024). To facilitate effective interaction, incorporating aspects of cognitive and emotional-social intelligence, the ability to recognize and interpret human emotions Mathur et al. (2024), can facilitate better interactions with more people. Factors such as age, cultural background, and personal experiences influence how individuals perceive and process information, particularly within subjective NLP tasks. Among these, emotion recognition (ER) and sentiment analysis (SA) are particularly sensitive to language- and culture-specific nuances Plaza-del-Arco et al. (2024).
@@ 3.1 Collecting Cultural Events
We manually craft scenarios, search on the web, and prompt LLMs to gather traditions, events, norms, and actions that elicit culturally different emotions across six target countries (UAE, USA, Germany, Ethiopia, India, and Mexico). We draw inspiration from the work of Havaldar et al. (2023b) to enhance topic diversity. Language representatives are asked to propose events distinct from those of other countries in the form of emotion-oriented questions. Importantly, these events do not contain explicit emotional keywords (as typically seen in traditional emotion datasets). We also refer to the International Survey on Emotion Antecedents and Reactions (ISEAR) (Scherer and Wallbott, 1994) data format, "When I … situations that cause a specific emotion", a well-known English dataset for emotion analysis consisting of self-reported events from around 3,000 respondents across 37 countries and five continents. Table 1 displays the ten broad categories of the CuLEmo dataset and the number of questions in each category.
@@
Table 1: The ten broad categories and the number of events (# Qn. - number of questions) in the CuLEmo dataset.
@@ 3.2 Human-Adapted Translation
After collecting the events in English, we translate them into five target languages—Arabic, Amharic, German, Hindi, and Spanish—using Google Translate, followed by native-speaker approvals. Because the questions are simple “How do you feel when … ?” questions and lack explicit emotion keywords, translation quality did not affect their cultural content. While we acknowledge that existing works often depend on translating emotion-annotated data from English into other languages with their labels Kajava et al. (2020); Tahir et al. (2023); Bianchi et al. (2022), our translation process is done before any annotation. To ensure correctness, native speakers evaluate the translations. While most translations were acceptable, a few adjustments were made, e.g., to fix gender references and timing expressions for Amharic.
@@ 3.4 CuLEmo Annotation
We use Amazon Mechanical Turk (MTurk) for most of our annotations, ensuring at least five native-speaker annotators per instance from the respective targeted country. We use a customized POTATO Pei et al. (2022) annotation tool for languages lacking sufficient MTurk annotators (e.g., Amharic) and recruit local native speakers who met our criteria. Annotators are fairly compensated \$12/hr (better than the Prolific [^3] https://www.prolific.com/’s minimum annotation wage, which is \$9/hr). Where no majority vote emerges among five annotators, we assign two additional annotations and use that majority vote.
@@ 5.1 Culture-Aware Emotion Prediction
Results: Table 2 presents the accuracy of each LLM for culture-aware emotion prediction. The choice of prompt language significantly influences performance. Generally, proprietary models are less affected by in-language prompts compared to open-source models, especially for Spanish and German. Certain cultures appear better represented in the models—Ministral-8B scores highly on German (72%), and GPT-4 performs best on Mexican (65%). In contrast, performance in Indian culture (Hindi) lags, particularly those using Hindi, Amharic, or Arabic scripts. Larger models do not always outperform smaller ones; Gemma-2-2B and Ministral-8B show competitive or superior accuracy relative to some proprietary models. When prompted in English, all models achieve a reasonable accuracy. Ministral-8B can exceed proprietary performance in English and German. GPT-4 answers the same emotion while using the corresponding country name in the prompting; see predicted examples in Table 3. Overall, results suggest that culture-aware emotion understanding remains challenging for the tested LLMs, especially for low-resource languages and cultures.